If f(x)=-3x+2,what are the values of m and b , so that , f(mx+b)=2x-3
Found 2 solutions by stanbon, ewatrrr:
Answer by stanbon(75887)

If f(x)=-3x+2,what are the values of m and b , so that , f(mx+b)=2x-3
---
f(mx+b) = -3(mx+b)+2
---
= -3mx - 3b+2
----
m = [-3m)
b = -3b+2
-----------------
Cheers,
Stan H.

Answer by ewatrrr(24785)

 
Hi,
f(x)=-3x+2
f(mx+b)= -3(mx+b)+2 = 2x-3
-3mx-3b = 2x-5
-3m = 2, m = -2/3
3b = 5, b = 5/3
CHECKING our Answer***
-3(mx+b)+2 = -3((-2/3)x -3(5/3)+2 = 2x -3